# Electricity Demand and Grid Impacts of AI Data Centers: Challenges and Prospects

### v3

@@ V-A2 Dynamics Modeling of AI Data Centers
High-fidelity modeling of the electrical dynamics of AI data center facilities is essential for assessing their grid disturbance ride-through capabilities and their interactions with power system stability. As inherently power electronics-based loads, AI data centers exhibit dynamic behavior governed by internal power converter controls and the temporal characteristics of computational workload processing. Hence, electromagnetic transient (EMT) modeling, using tools such as EMTP or PSCAD, is required for capturing the fast, high-frequency responses of these AI facilities, which may not be accurately represented by conventional phasor-domain models [121, 122]. A recent dynamic load model tailored for AI data centers is proposed in [123] that incorporates three key features, including an upstream UPS, cooling loads represented by induction motor dynamics, and pulsing compute loads to represent AI workload transients. Complementary approaches include hierarchical and hybrid modeling frameworks that combine high-fidelity EMT simulations for critical components with reduced-order equivalents for external network regions. This method balances computational efficiency with accuracy, enabling reliable simulation of real-world grid impacts from AI data center behavior [124, 125].
@@ V-A3 Tailored Demand Response Program Design
To effectively harness the flexibility of AI data centers while minimizing adverse operational impacts, grid operators need to design demand response (DR) programs specifically tailored to the unique characteristics of these loads. Standard price-based DR mechanisms, while widely used in residential and industrial contexts, may introduce new challenges when applied directly to large-scale AI loads. For instance, inference or training tasks deferred in response to high electricity prices may be rescheduled simultaneously once prices fall, leading to secondary demand peaks or rebound effects that stress the grid. Moreover, AI data center operators are relatively insensitive to electricity prices. As noted in [126], these facilities have high sunk capital costs, and the workloads they run are highly lucrative. As a result, the opportunity cost of deferring training or inference tasks can outweigh potential savings from energy arbitrage. Therefore, relying on price signals alone may be insufficient to drive effective load shifting. This underscores the need for carefully designed DR mechanisms tailored to AI data centers. One promising direction involves incentive-based (or penalty-based) and contract-based DR mechanisms, in which AI data centers commit to predefined levels of flexibility in exchange for direct compensation or priority grid services. For example, Google has signed contractual agreements with utilities in Indiana and Tennessee to curtail AI workloads during periods of grid stress [127, 128]. Additionally, Coincident Peak (CP) mechanisms, such as ERCOT’s Four Coincident Peak (4CP) charge, which imposes significant penalties on energy consumption during the grid’s four highest monthly demand intervals, can directly incentivize data centers to shift their usage away from peak periods [129].
@@ V-A4 Standardization and Regulation
As the scale and concentration of AI data centers increase, the implementation of specific standards and regulatory frameworks becomes essential to confine their load behavior and preserve grid reliability. One key task is the specification of fault ride-through (FRT) capabilities. Similar to requirements imposed on renewable energy sources under IEEE Standard 1547 and European grid codes, AI data centers may be required to remain connected during short-term voltage or frequency disturbances. For instance, proposed specifications for data center UPS systems suggest maintaining operation down to 50–70% of nominal voltage with a resynchronization capability within one second [130, 131]. In addition to ride-through, dynamic performance requirements should address acceptable voltage and frequency tolerance ranges, reactive power support, harmonic emissions, and ramp rate limits. Such performance standards are increasingly being applied to large electrical and hybrid loads under evolving European interconnection frameworks [132]. Moreover, interconnection standards may require that AI data center operators provide utilities with accurate dynamic models, anonymized workload profiles, and operational telemetry to enable realistic system planning and stability analysis [133]. On the regulatory front, the North American Electric Reliability Corporation (NERC) has launched a Large Loads Task Force to assess and manage risks associated with high-density data center growth [134]. There is growing momentum for applying grid codes to large, sensitive electronic loads, including mandatory response characteristics during contingencies [135]. As a result, these regulatory and standardization measures are critical to ensuring that the integration of AI data centers supports, rather than undermines, the stability and resilience of the power grid.
@@ V-B1 Leveraging Temporal and Spatial Flexibility of AI Data Centers for Grid-Aware Operation
Flexible AI workloads can be rescheduled or shifted without compromising service quality, offering valuable flexibility to support grid operation. Specifically, temporal flexibility refers to the ability to shift AI workloads over time without compromising service quality or user experience. For example, non-urgent computational tasks, such as large-scale model training or batch inference, can be scheduled during off-peak hours or periods of low grid stress and high renewable energy availability. This helps to flatten load curves, enhance grid utilization, lower operating costs, and reduce carbon emissions. Spatial flexibility refers to the capability to reroute AI workloads across multiple data centers in different geographic locations. This allows AI workloads to be rerouted to facilities in regions with cleaner electricity, lower prices, or less grid congestion. Together, temporal and spatial flexibility offer powerful levers for optimizing the alignment between AI demand and power system conditions, helping to enhance grid reliability, reduce carbon emissions, and lower operational costs. Prior research shows that temporally shifting AI workloads can reduce operational costs by up to 12% and carbon emissions by approximately 10% [136, 137, 138]. Real-world implementations [139] are increasingly supported by energy-aware scheduling platforms and demand-sensing technologies [140]. Moreover, studies based on locational marginal emissions modeling highlight that even modest spatial shifting of loads can yield significant emission reductions [141]. In addition, AI data centers equipped with advanced control systems and on-site energy storage can mitigate their own power fluctuations [18] and even provide ancillary services to the grid. By modulating compute intensity or charging and discharging stored energy in response to grid signals, such facilities can contribute to frequency regulation, spinning reserves, and fast-ramping support [142, 14]. This positions AI data centers not only as large load consumers but also as active grid participants that enhance grid reliability [143].
V-B2 Energy-Aware AI Workload and Hardware
@@
(i) Energy-aware AI computing: Techniques such as energy-aware training (EAT) [144] enable models to activate only the neurons and connections necessary for a given task, effectively disabling unused components and lowering power use. Additional savings can be achieved through dynamic voltage and frequency scaling of CPUs and GPUs. For example, reference [126] reports that under-clocking a single NVIDIA A100 GPU for a specific application reduced power consumption by 40% while decreasing performance by only 22%. This demonstrates that certain jobs could be executed at lower frequencies over a longer duration to reduce overall energy use. Moreover, advanced power smoothing and management techniques at the software and GPU levels can be developed and deployed to mitigate power swings induced by AI workloads [18], thus reducing risks to grid stability.
@@
(ii) Development of lighter, task-specific AI models: Smaller, task-specific AI models can match the accuracy of large foundation models while consuming substantially less power. Techniques such as pruning (removing redundant parameters), quantization (reducing numerical precision), and knowledge distillation (transferring knowledge from a large model to a smaller one) lower computational demands without compromising output quality. These approaches also facilitate deployment on edge and mobile devices while yielding significant energy savings. For example, TinyBERT achieves over 90% lower energy consumption compared to full-size BERT models [145]. Furthermore, sparse model architectures enhance efficiency by directing computational resources toward the most critical regions of the network.
@@
(iii) Energy-efficient hardware design: The deployment of energy-efficient hardware are necessary for reducing the carbon and energy footprint of AI data centers. Advances in processor architecture can substantially improve performance-per-watt, enabling higher computational throughput at lower power costs. Domain-specific accelerators, low-power CPUs/GPUs, and AI-optimized hardware such as Google’s TPUs are central to this effort. For instance, Google’s TPU v4-based supercomputer has demonstrated notable speed and efficiency gains over NVIDIA A100 GPUs [146]. Moreover, AI itself can be leveraged to optimize the performance and energy efficiency of its supporting infrastructure, as illustrated in [147]. Co-optimizing AI model architectures with underlying hardware, combining techniques such as pruning, quantization, and hardware-aware model design, offers further potential for achieving substantial energy savings without sacrificing accuracy.
V-B3 Smart and Efficient Cooling Techniques
@@
Cooling facilities account for the second-largest share of electricity consumption in AI data centers. Recent research and industry trends emphasize several advanced strategies: (i) Hybrid air-liquid cooling systems: combining air-based airflow management (e.g., hot/cold aisle containment, in-row cooling, raised-floor systems) with liquid approaches (e.g., direct-to-chip) to optimize energy, complexity, and adaptability. These systems help manage thermal loads from dense AI clusters while reducing energy usage compared to air-only or liquid-only designs [148, 149]. (ii) Immersion and direct liquid cooling: immersing servers in dielectric fluids or using piping/driven liquid heat removal offers dramatically improved heat transfer, higher rack density, and up to 50% energy savings compared to air cooling, though challenges remain around maintenance and cost [150, 151]. (iii) AI-driven adaptive and predictive cooling control: leveraging reinforcement learning (RL) [152, 153, 154], deep neural networks, or other ML models to anticipate thermal load changes and dynamically tune cooling systems for maximum efficiency. These approaches have achieved 14-21% energy savings in live data center deployments without violating safety constraints [155]. Google’s commercial cooling experiments using RL yielded 9-13% energy reductions [156]. (iv) Renewable-powered and free (economizer) cooling: leveraging ambient air (free cooling) where climates permit, and powering cooling infrastructure with onsite renewables (solar, wind) to reduce carbon intensity. These approaches support sustainable operations while attenuating grid demand [157]. These smart and efficient cooling strategies underscore a shift towards multi-dimensional optimization to balance energy, water, carbon, and operational complexity.
@@ V-C1 Energy-Aware Prompting and AI Model Selection
End-users and developers of AI applications can contribute to reducing the AI electricity footprint through energy-aware application development and informed usage decisions. For end-users, one critical approach is the adoption of energy-efficient prompting strategies and the selection of appropriately scaled models. Recent studies have shown that longer, more verbose prompts, while potentially improving model interpretability or output alignment, can significantly increase the number of tokens to be processed and thus greatly elevate the power consumption of LLMs [158, 159, 160]. Moreover, fine-tuned, task-specific lightweight AI models often yield comparable performance to general-purpose foundation models but at a small fraction of the computational and energy cost [161]. Recent initiatives, such as the AI Energy Score project [162], aim to bring transparency to model energy efficiency by benchmarking the energy cost per query across different AI models. Making such metrics available to end-users can encourage more responsible usage behaviors and incentivize developers to adopt energy-aware design considerations.
@@ V-C2 User Flexibility and AI Demand Response
In practice, many end-users of AI applications exhibit a degree of flexibility in their usage behaviors, as they often do not need immediate AI responses to their queries. For example, a user submitting a complex AI task late at night may not need the results until the following morning. This inherent temporal flexibility in AI demand of end-users offers a valuable opportunity to better align query processing with power grid operational needs. Building on this insight, a novel concept, termed “AI Demand Response (AI-DR)”, can be developed to harness user-side AI workload flexibility, analogous to the electric demand response programs [161, 163, 164] in power systems. Similarly, there can be incentive-based and price-based AI-DR mechanisms. In incentive-based AI-DR programs, end-users are incentivized through lower subscription fees, discounted inference rates, or access to premium model tiers in exchange for accepting minor service compromises [126]. These may include bounded response delays, use of lower-tier hardware, or allowing their queries to be processed during periods of low grid stress or high renewable generation. In contrast, price-based AI-DR programs employ time-varying pricing schemes, such as real-time pricing, time-of-use, critical peak pricing [161], to encourage users to shift their AI usage away from peak hours. By exposing end-users to dynamic AI usage pricing signals, service providers can influence AI workload patterns in a grid-aware and energy-efficient manner. When aggregating across a large and diverse user base, service providers (AI data centers) can provide substantial flexibility to support grid reliability and improve overall cost efficiency.

### v4

@@ V-A2 Dynamics Modeling of AI Data Centers
High-fidelity modeling of the electrical dynamics of AI data center facilities is essential for assessing their grid disturbance ride-through capabilities and their interactions with power system stability. As inherently power electronics-based loads, AI data centers exhibit dynamic behavior governed by internal power converter controls and the temporal characteristics of computational workload processing. Hence, electromagnetic transient (EMT) modeling, using tools such as EMTP or PSCAD, is required for capturing the fast, high-frequency responses of these AI facilities, which may not be accurately represented by conventional phasor-domain models [allabadi2024emt, search11_largeLoads]. A recent dynamic load model tailored for AI data centers is proposed in [jimenez2025dynamic] that incorporates three key features, including an upstream UPS, cooling loads represented by induction motor dynamics, and pulsing compute loads to represent AI workload transients. Complementary approaches include hierarchical and hybrid modeling frameworks that combine high-fidelity EMT simulations for critical components with reduced-order equivalents for external network regions. This method balances computational efficiency with accuracy, enabling reliable simulation of real-world grid impacts from AI data center behavior [search25_dynReduction, academia23_hmm].
@@ V-A3 Tailored Demand Response Program Design
To effectively harness the flexibility of AI data centers while minimizing adverse operational impacts, grid operators need to design demand response (DR) programs specifically tailored to the unique characteristics of these loads. Standard price-based DR mechanisms, while widely used in residential and industrial contexts, may introduce new challenges when applied directly to large-scale AI loads. For instance, inference or training tasks deferred in response to high electricity prices may be rescheduled simultaneously once prices fall, leading to secondary demand peaks or rebound effects that stress the grid. Moreover, AI data center operators are relatively insensitive to electricity prices. As noted in [crozier_potential_2025], these facilities have high sunk capital costs, and the workloads they run are highly lucrative. As a result, the opportunity cost of deferring training or inference tasks can outweigh potential savings from energy arbitrage. Therefore, relying on price signals alone may be insufficient to drive effective load shifting. This underscores the need for carefully designed DR mechanisms tailored to AI data centers. One promising direction involves incentive-based (or penalty-based) and contract-based DR mechanisms, in which AI data centers commit to predefined levels of flexibility in exchange for direct compensation or priority grid services. For example, Google has signed contractual agreements with utilities in Indiana and Tennessee to curtail AI workloads during periods of grid stress [turn0news11, turn0news10]. Additionally, Coincident Peak (CP) mechanisms, such as ERCOT’s Four Coincident Peak (4CP) charge, which imposes significant penalties on energy consumption during the grid’s four highest monthly demand intervals, can directly incentivize data centers to shift their usage away from peak periods [Lukawski2019DemandResponse].
@@ V-A4 Standardization and Regulation
As the scale and concentration of AI data centers increase, the implementation of specific standards and regulatory frameworks becomes essential to confine their load behavior and preserve grid reliability. One key task is the specification of fault ride-through (FRT) capabilities. Similar to requirements imposed on renewable energy sources under IEEE Standard 1547 and European grid codes, AI data centers may be required to remain connected during short-term voltage or frequency disturbances. For instance, proposed specifications for data center UPS systems suggest maintaining operation down to 50–70% of nominal voltage with a resynchronization capability within one second [turn0search3, turn0search20]. In addition to ride-through, dynamic performance requirements should address acceptable voltage and frequency tolerance ranges, reactive power support, harmonic emissions, and ramp rate limits. Such performance standards are increasingly being applied to large electrical and hybrid loads under evolving European interconnection frameworks [turn0search12]. Moreover, interconnection standards may require that AI data center operators provide utilities with accurate dynamic models, anonymized workload profiles, and operational telemetry to enable realistic system planning and stability analysis [gridlab2025]. On the regulatory front, the North American Electric Reliability Corporation (NERC) has launched a Large Loads Task Force to assess and manage risks associated with high-density data center growth [turn0search2]. There is growing momentum for applying grid codes to large, sensitive electronic loads, including mandatory response characteristics during contingencies [turn0search8]. As a result, these regulatory and standardization measures are critical to ensuring that the integration of AI data centers supports, rather than undermines, the stability and resilience of the power grid.
@@ V-B1 Leveraging Temporal and Spatial Flexibility of AI Data Centers for Grid-Aware Operation
Flexible AI workloads can be rescheduled or shifted without compromising service quality, offering valuable flexibility to support grid operation. Specifically, temporal flexibility refers to the ability to shift AI workloads over time without compromising service quality or user experience. For example, non-urgent computational tasks, such as large-scale model training or batch inference, can be scheduled during off-peak hours or periods of low grid stress and high renewable energy availability. This helps to flatten load curves, enhance grid utilization, lower operating costs, and reduce carbon emissions. Spatial flexibility refers to the capability to reroute AI workloads across multiple data centers in different geographic locations. This allows AI workloads to be rerouted to facilities in regions with cleaner electricity, lower prices, or less grid congestion. Together, temporal and spatial flexibility offer powerful levers for optimizing the alignment between AI demand and power system conditions, helping to enhance grid reliability, reduce carbon emissions, and lower operational costs. Prior research shows that temporally shifting AI workloads can reduce operational costs by up to 12% and carbon emissions by approximately 10% [frontiers2024temporal, sciencedirect2024spatial, nature2024spatial]. Real-world implementations [radovanovic2022carbon] are increasingly supported by energy-aware scheduling platforms and demand-sensing technologies [wiki_greencomputing]. Moreover, studies based on locational marginal emissions modeling highlight that even modest spatial shifting of loads can yield significant emission reductions [arxiv2020marginalemission]. In addition, AI data centers equipped with advanced control systems and on-site energy storage can mitigate their own power fluctuations [choukse2025power] and even provide ancillary services to the grid. By modulating compute intensity or charging and discharging stored energy in response to grid signals, such facilities can contribute to frequency regulation, spinning reserves, and fast-ramping support [ela2011assessment, aljbour2024powering]. This positions AI data centers not only as large load consumers but also as active grid participants that enhance grid reliability [spieler_how_2025].
V-B2 Hybrid Energy Storage Solutions
Energy storage systems are increasingly essential for mitigating the fast and large power fluctuations of AI data centers. However, individual energy storage technologies often fall short of meeting the full spectrum of performance requirements, including rapid response, adequate power and energy capacity, long cycle life, and cost-effectiveness. As a result, hybrid energy storage systems (HESS), which integrate multiple storage technologies with complementary operational characteristics, have emerged as a promising solution to address these diverse demands [wang2012energy]. Specifically, supercapacitors [zheng2016hybrid, 10839152] offer sub-second response times and exceptionally high cycle life, making them well-suited for absorbing ultra-short-term power spikes at the server or rack level. Batteries, particularly lithium-ion and emerging alternatives such as sodium-ion, provide higher energy density and longer-duration buffering, which makes them suitable for load shifting, peak shaving, and backup at the facility level [guo2021integrated]. Flywheels [amiryar2017review] deliver medium-duration, high-power buffering with high round-trip efficiency, making them effective for smoothing aggregated power fluctuations and providing short-term ride-through support at the cluster level [arghandeh2012flywheel]. A potential hybrid configuration, in which supercapacitors manage millisecond-to-second transients, flywheels address second-to-minute ramps, and batteries handle minute-to-hour variations, offers a balance of technical performance, lifecycle efficiency, and overall cost-effectiveness. As AI workloads continue to grow in intensity and variability, the deployment of hybrid storage architectures, coupled with smart energy storage management systems, will play a critical role in enabling sustainable, reliable, and grid-friendly AI data center operations.
V-B3 Energy-Aware AI Workload and Hardware
@@
(i) Energy-aware AI computing: Techniques such as energy-aware training (EAT) [lazzaro2023minimizing] enable models to activate only the neurons and connections necessary for a given task, effectively disabling unused components and lowering power use. Additional savings can be achieved through dynamic voltage and frequency scaling of CPUs and GPUs. For example, reference [crozier_potential_2025] reports that under-clocking a single NVIDIA A100 GPU for a specific application reduced power consumption by 40% while decreasing performance by only 22%. This demonstrates that certain jobs could be executed at lower frequencies over a longer duration to reduce overall energy use. Moreover, advanced power smoothing and management techniques at the software and GPU levels can be developed and deployed to mitigate power swings induced by AI workloads [choukse2025power], thus reducing risks to grid stability.
@@
(ii) Development of lighter, task-specific AI models: Smaller, task-specific AI models can match the accuracy of large foundation models while consuming substantially less power. Techniques such as pruning (removing redundant parameters), quantization (reducing numerical precision), and knowledge distillation (transferring knowledge from a large model to a smaller one) lower computational demands without compromising output quality. These approaches also facilitate deployment on edge and mobile devices while yielding significant energy savings. For example, TinyBERT achieves over 90% lower energy consumption compared to full-size BERT models [paula2025comparative]. Furthermore, sparse model architectures enhance efficiency by directing computational resources toward the most critical regions of the network.
@@
(iii) Energy-efficient hardware design: The deployment of energy-efficient hardware are necessary for reducing the carbon and energy footprint of AI data centers. Advances in processor architecture can substantially improve performance-per-watt, enabling higher computational throughput at lower power costs. Domain-specific accelerators, low-power CPUs/GPUs, and AI-optimized hardware such as Google’s TPUs are central to this effort. For instance, Google’s TPU v4-based supercomputer has demonstrated notable speed and efficiency gains over NVIDIA A100 GPUs [google2023tpuv4]. Moreover, AI itself can be leveraged to optimize the performance and energy efficiency of its supporting infrastructure, as illustrated in [bolon2024green]. Co-optimizing AI model architectures with underlying hardware, combining techniques such as pruning, quantization, and hardware-aware model design, offers further potential for achieving substantial energy savings without sacrificing accuracy.
V-B4 Smart and Efficient Cooling Techniques
@@
Cooling facilities account for the second-largest share of electricity consumption in AI data centers. Recent research and industry trends emphasize several advanced strategies: (i) Hybrid air-liquid cooling systems: combining air-based airflow management (e.g., hot/cold aisle containment, in-row cooling, raised-floor systems) with liquid approaches (e.g., direct-to-chip) to optimize energy, complexity, and adaptability. These systems help manage thermal loads from dense AI clusters while reducing energy usage compared to air-only or liquid-only designs [TakingHybrid2024, GuideToCooling2025]. (ii) Immersion and direct liquid cooling: immersing servers in dielectric fluids or using piping/driven liquid heat removal offers dramatically improved heat transfer, higher rack density, and up to 50% energy savings compared to air cooling, though challenges remain around maintenance and cost [ImmersionCooling2025, mohammedPerformanceImprovementHighdensity2025]. (iii) AI-driven adaptive and predictive cooling control: leveraging reinforcement learning (RL) [chen2022reinforcement, kahil2025reinforcement, ran2022optimizing], deep neural networks, or other ML models to anticipate thermal load changes and dynamically tune cooling systems for maximum efficiency. These approaches have achieved 14-21% energy savings in live data center deployments without violating safety constraints [Zhan2025DC21]. Google’s commercial cooling experiments using RL yielded 9-13% energy reductions [Luo2022RL]. (iv) Renewable-powered and free (economizer) cooling: leveraging ambient air (free cooling) where climates permit, and powering cooling infrastructure with onsite renewables (solar, wind) to reduce carbon intensity. These approaches support sustainable operations while attenuating grid demand [oro2015energy]. These smart and efficient cooling strategies underscore a shift towards multi-dimensional optimization to balance energy, water, carbon, and operational complexity.
@@ V-C1 Energy-Aware Prompting and AI Model Selection
End-users and developers of AI applications can contribute to reducing the AI electricity footprint through energy-aware application development and informed usage decisions. For end-users, one critical approach is the adoption of energy-efficient prompting strategies and the selection of appropriately scaled models. Recent studies have shown that longer, more verbose prompts, while potentially improving model interpretability or output alignment, can significantly increase the number of tokens to be processed and thus greatly elevate the power consumption of LLMs [rubei2025prompt, adamska2025green, husom2024price]. Moreover, fine-tuned, task-specific lightweight AI models often yield comparable performance to general-purpose foundation models but at a small fraction of the computational and energy cost [OCONNELL2014686]. Recent initiatives, such as the AI Energy Score project [turn0search14], aim to bring transparency to model energy efficiency by benchmarking the energy cost per query across different AI models. Making such metrics available to end-users can encourage more responsible usage behaviors and incentivize developers to adopt energy-aware design considerations.
@@ V-C2 User Flexibility and AI Demand Response
In practice, many end-users of AI applications exhibit a degree of flexibility in their usage behaviors, as they often do not need immediate AI responses to their queries. For example, a user submitting a complex AI task late at night may not need the results until the following morning. This inherent temporal flexibility in AI demand of end-users offers a valuable opportunity to better align query processing with power grid operational needs. Building on this insight, a novel concept, termed “AI Demand Response (AI-DR)”, can be developed to harness user-side AI workload flexibility, analogous to the electric demand response programs [OCONNELL2014686, jordehi2019optimisation, chen2021online] in power systems. Similarly, there can be incentive-based and price-based AI-DR mechanisms. In incentive-based AI-DR programs, end-users are incentivized through lower subscription fees, discounted inference rates, or access to premium model tiers in exchange for accepting minor service compromises [crozier_potential_2025]. These may include bounded response delays, use of lower-tier hardware, or allowing their queries to be processed during periods of low grid stress or high renewable generation. In contrast, price-based AI-DR programs employ time-varying pricing schemes, such as real-time pricing, time-of-use, critical peak pricing [OCONNELL2014686], to encourage users to shift their AI usage away from peak hours. By exposing end-users to dynamic AI usage pricing signals, service providers can influence AI workload patterns in a grid-aware and energy-efficient manner. When aggregating across a large and diverse user base, service providers (AI data centers) can provide substantial flexibility to support grid reliability and improve overall cost efficiency.